OpenStudy (anonymous):

Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square

OpenStudy (p0sitr0n):

f(x) = -2(x-5)^2+5 its the canonical form

OpenStudy (p0sitr0n):

graphine = enter on calculator and find the intersection with the axis x

OpenStudy (p0sitr0n):

\[x1,x2 = (-b \pm \sqrt{b^2-4ac})/2a\] a is -2, b = 20 c =-45 solve normal formula of the same equation is -2x^2 +20x -45

OpenStudy (p0sitr0n):

square completion -2x^2 +20x -45=0 x^2 -10x +45/2 = 0 continue